OpenStudy (anonymous):

find the slope (2,-1) (6,3)

OpenStudy (aroub):

one

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[slope=(y2-y1)/(x2-x1)\]

OpenStudy (anonymous):

3--1 / 6-2 = 4/4 = 1